# Adiabatic theorem for non-Hermitian quantum systems with non-degenerate real eigenvalues: A proof following Kato’s approach

### v1

@@ IV Discussions and Conclusion
Apparently, the result is somewhtat intuitive. One may wonder why such a result deserves a specified proof. In fact, any Hamiltonian $H(s)$ with non-degenerate real eigenvalues is similar to a Hermitian Hamiltonian $H_{2}(s)$. If $H(s)=V^{-1}H_{2}(s)V$, then the evolution operator of $U_{T}(s)=V^{-1}U^{(2)}_{T}(s)$, where $U^{(2)}_{T}(s)$ is the evolution operator of $TH_{2}(s)$. Moreover, if $x(s)$ is an eigenvector of $H_{2}(s)$, $V^{-1}x(s)$ is an eigenvector of $H(s)$. Now since the adiabatic theorem is valid for $H_{2}(s)$, it is also valid for $H(s)$. However, in general we have $H(s)=V^{-1}(s)H_{2}(s)V(s)$, the matrix $V(s)$ is not independent of $s$, thus $U_{T}(s)\neq V^{-1}(s)U^{(2)}_{T}(s)$. In this case, one cannot directly obtain the adiabatic theorem for $H(s)$ from $H_{2}(s)$. Hence a proof is necessary.
Unlike the Hermitian case, the eigenvectors are usually unnormalized in the non-Hermitian case. Nevertheless, multiplying the eigenvector by some factor will bring no essential change to the result. In fact, we consider a matrix $\tilde{V}(s)$, whose column vectors are $v_{j}(s)e^{ir_{j}(s)}$, where $r_{j}(s)=\int_{0}^{s}i\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma$. In other words, the matrix $\tilde{V}(s)$ only depends on the initial value $V(0)$ and is independent of the choice of the eigenvectors.
Suppose that we have another choice of eigenvectors $\psi_{j}(s)$, and $\phi_{i}(s)$ form a biorthogonal system, i.e. $\phi^{\dagger}_{i}(s)\psi_{j}(s)=\delta_{ij}$. Then it is obvious that $\psi_{j}(s)=\mu_{j}(s)v_{j}(s)$ and $\phi_{j}(s)=\frac{1}{\overline{\mu}_{j}(s)}\xi_{j}(s)$, where $\mu_{j}(s)$ is some complex number and $\overline{\mu}_{j}(s)$ is its conjugate.
@@
Now suppose that one has an initial state (unnormalized vector) $q_{j}(0)=c_{j}(0)v_{j}(0)$. The above derivation and direct calculations show that the adiabatically evolved state $q_{j}(s)$ can always be written as $q_{j}(s)=c_{j}(0)v_{j}(s)e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}$ regardless of using $v_{j}$ or $\psi_{j}$ as the reference in calculation. It should also be mentioned that the Berry phase in our discussion is complex, which is similar to [21].
Even though our proof is inspired by Kato’s, there are also differences. The first is Kato’s work only applies to Hermitian systems since the utilizing the properties of orthogonal projections and unitary matrix. However, the result in this paper also applies to the non-Hermitian case. Moreover, Kato’s work does not involve the concept of Berry phase, while the Berry phase in our discussion plays an important role. In fact, Kato’s work discussed the solution to the differential equation of the orthogonal projections. This procedure is not needed in this paper, such a simplification is due to the use of Berry phase. There is one place where this paper is more complex than Kato’s work. We have to show the uniform boundedness of $U_{T}(s)$ and $U_{T}^{-1}(s)$, while in Kato’s discussion on Hermitian systems, $U_{T}(s)$ is unitary and is obvious uniformly bounded.
In conclusion, our result generalizes the adiabatic theorem, which is still valid for non-Hermitian systems with non-degenerate real eigenvalues.
@@ V Appendix
V.1 The Berry phase
The concrete concept about Berry phase appears much later than the adiabatic theorem. Nowadays, in the discussions of adiabatic theorem, such a Berry phase term is often taken into account. A question naturally arises on how Kato’s approach in proving the adiabatic theorem did not utilize the Berry phase.
In fact, in such a discussion, one may choose a gauge such that the Berry phase can be cancelled out. Thus the adiabatic theorem can still be valid. In fact, Kato’s discussion of the differential equation of the projection plays such a role. A similar discussion can also be found in [20]. Of course, when the Hamiltonian is cyclic, the Berry phase cannot omitted in general. Hence it is better to keep it.
To see this, note that if
$U_{T}(s)\phi_{j}(0)=e^{-iT\nu_{j}(s)+ir_{j}(s)}\phi_{j}(s),$
then one can take $\psi_{j}(s)=\alpha_{j}(s)\phi_{j}(s)$, where $\alpha_{j}(s)=\alpha_{j}(0)e^{ir_{j}(s)}$ and $\alpha_{j}(0)$ is some constant. It is obvious that $U_{T}(s)\psi_{j}(0)=e^{-iT\nu_{j}(s)}\psi_{j}(s)$. That is, the Berry phase can be absorbed. However, when $H(s^{\prime})=H(0)$, the above calculation does not guarantee that $\phi_{j}(s^{\prime})=\phi_{j}(0)$. If we add the condition that $\phi_{j}(s^{\prime})=\phi_{j}(0)$, then $U_{T}(s^{\prime})\phi_{j}(0)=e^{-iT\nu_{j}(s^{\prime})+ir_{j}(s^{\prime})}\phi_{j}(0)$. Moreover, it is demanded that $\psi_{j}(s^{\prime})=\psi_{j}(0)$. Now direct calculations show that
$U_{T}(s^{\prime})\psi_{j}(0)=e^{-iT\nu_{j}(s)+ir_{j}(s)}\psi_{j}(0)=e^{-iT\nu_{j}(s)+ir_{j}(s)}\psi_{j}(s^{\prime}).$
Thus, the Berry phase cannot be omitted in the cyclic case. Put it another way, if we simply write the $U_{T}(s)\psi_{j}(0)=e^{-iT\nu_{j}(s)}\psi_{j}(s)$, even when $H(s^{\prime})=H(0)$, $\psi_{j}(s^{\prime})\neq\psi_{j}(0)$. They differ in a factor. However, in general it is demanded that $\psi_{j}(s^{\prime})=\psi_{j}(0)$. In this case, the Berry phase cannot be omitted. In Refs. [1, 20], the Berry phase is not explicitly written out, but adsorbed with the adiabatic theorem.
V.2 The Grönwall inequality
@@ Acknowledgement
This work is partially supported by the National Natural Science Foundation of China (12371135). The first author thank Prof. Dianmin Tong and Zeqian Chen for useful suggestions.

### v2

@@ IV Discussions and Conclusion
The above discussion shows that the adiabatic theorem is valid for the diagonalizable Hamiltonians with real eigenvalues. That is, the eigenstate will remain in the subspace when evolving slowly enough. When the energy is non-degenerate, it reduces to the Berry phase.
Apparently, the result is somewhat natural. One may wonder why such a result deserves a specified proof. In fact, any Hamiltonian $H(s)$ with non-degenerate real eigenvalues is similar to a Hermitian Hamiltonian $H_{2}(s)$. If $H(s)=V^{-1}H_{2}(s)V$, then the evolution operator of $U_{T}(s)=V^{-1}U^{(2)}_{T}(s)$, where $U^{(2)}_{T}(s)$ is the evolution operator of $TH_{2}(s)$. Moreover, if $x(s)$ is an eigenvector of $H_{2}(s)$, then $V^{-1}x(s)$ is an eigenvector of $H(s)$. Now since $H_{2}(s)$ is Hermitian, the adiabatic theorem is valid for $H_{2}(s)$. Moreover, since $V$ is time independent and $U_{T}(s)=V^{-1}U^{(2)}_{T}(s)$, one can directly verify that the adiabatic theorem is also valid for $H(s)$. However, when $H(s)=V^{-1}(s)H_{2}(s)V(s)$, i.e. the matrix $V(s)$ is dependent of $s$, one has $U_{T}(s)\neq V^{-1}(s)U^{(2)}_{T}(s)$. In this case, one cannot directly obtain the adiabatic theorem for $H(s)$ from $H_{2}(s)$. Hence a proof is necessary.
Unlike the Hermitian case, the eigenvectors of the Hamiltonian $H(s)$ are usually unnormalized in the non-Hermitian case for convenience. That is, the eigenvectors are not uniquely defined since their norms are undetermined. Consequently, the Berry phases are not uniquely determined either. However, this will bring no essential change to the result. To see this, let $v_{j}(s)$ be the eigenvectors of $H(s)$ and $\xi_{j}(s)$ form a biorthogonal system. Assume that there is another choice of eigenvectors $\psi_{j}(s)$ and $\phi_{i}(s)$ form a biorthogonal system. Now since $H(s)$ has nondegenerate eigenvalues, it is obvious that $\psi_{j}(s)=\mu_{j}(s)v_{j}(s)$ and $\phi_{j}(s)=\frac{1}{\overline{\mu}_{j}(s)}\xi_{j}(s)$, where $\mu_{j}(s)$ is some complex number and $\overline{\mu}_{j}(s)$ is its conjugate.
@@
Now suppose that one has an initial state (unnormalized vector) $q_{j}(0)=c_{j}(0)v_{j}(0)$. The above derivation and direct calculations show that the adiabatically evolved state $q_{j}(s)$ can always be written as $q_{j}(s)=c_{j}(0)v_{j}(s)e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}$ regardless of using $v_{j}$ or $\psi_{j}$ as the reference in calculation. It should also be mentioned that the Berry phase in our discussion is complex, which is similar to [21]. In fact, a complex Berry phase and the biorthogonal states are also used in [30].
In Eqs. (6)-(8), our proof utilizes the spectral decomposition and functional calculus with respect to biorthogonal systems. Such an approach is inspired by Kato’s discussions on Hermitian systems, which can avoid the Riemann-Lebesgue Lemma in the discussion [2]. However, since its utilization of orthogonal projections and unitary matrices, Kato’s discussions do not directly apply to the non-Hermitian case. Moreover, without the Berry phase or its non-abelian counterpart, a treatment of the differential equation of the orthogonal projections and the intertwining operator is necessary in the proof. This long procedure is not needed in this paper. Due to the use of Berry phase or the non-abelian geometric phase, the proof is largely simplified.
There is one place where this paper is more complex than the Hermitian case. One has to show the uniform boundedness of $U_{T}(s)$ and $U_{T}^{-1}(s)$, which are obvious in Hermitian quantum mechanics. To show the uniform boundedness, we use the Grönwall inequality. Given that our discussion are in a more general case, this complication is not unexpected.
In conclusion, we obtained a more general adiabatic theorem for non-Hermitian systems with real eigenvalues. The discussions also partly justifies the definition of a complex Berry phase and its non-abelian generalization.
@@ V Appendix
V.1 The Grönwall inequality
@@ Acknowledgement
This work is partially supported by the National Natural Science Foundation of China (12371135). The first author thanks Prof. Chunlan Jiang, Dianmin Tong and Zeqian Chen for useful suggestions.